### v2

# Efficiently and Effectively: A Two-stage Approach to Balance Plaintext and Encrypted Text for Traffic Classification

Abstract.
Encrypted traffic classification is the task of identifying the application or service associated with encrypted network traffic. One effective approach for this task is to use deep learning methods to encode the raw traffic bytes directly and automatically extract features for classification (byte-based models). However, current byte-based models input raw traffic bytes, whether plaintext or encrypted text, for automated feature extraction, neglecting the distinct impacts of plaintext and encrypted text on downstream tasks. Additionally, these models primarily focus on improving classification accuracy, with little emphasis on the efficiency of models. In this paper, for the first time, we analyze the impact of plaintext and encrypted text on the model’s effectiveness and efficiency. Based on our observations and findings, we propose a two-phase approach to balance the trade-off between plaintext and encrypted text in traffic classification. Specifically, Stage one is to Determine whether the Plain text is enough to be accurately Classified (DPC) using the proposed DPC Selector. This stage quickly identifies samples that can be classified using plaintext, leveraging explicit byte features in plaintext to enhance model’s efficiency. Stage two aims to adaptively make a classification with the result from stage one. This stage incorporates encrypted text information for samples that cannot be classified using plaintext alone, ensuring the model’s effectiveness on traffic classification tasks. Experiments on two datasets demonstrate that our proposed model achieves state-of-the-art results in both effectiveness and efficiency.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Encrypted traffic classification is a critical task in the field of cybersecurity, focusing on identifying and categorizing network traffic data that has been encrypted (Bujlow et al., 2015; Aceto et al., 2020). This process is essential for maintaining network security, managing bandwidth, detecting malicious activities, and ensuring regulatory compliance (Akbari et al., 2021; Cui et al., 2023). Applications of encrypted traffic classification span various domains, including intrusion detection systems, quality of service (QoS) management, and network forensics. As illustrated in Figure 1, some common applications include encrypted application classification (Lin et al., 2022), attack detection (Meng et al., 2021), VPN classification (Draper-Gil et al., 2016), etc.
Figure 1. An example of traffic data packet. The right-hand part means the application classification task. Green font represents the different categories.
Figure 2. Testing time and accuracy analysis on different settings between plaintest and encrypted text on four datasets.
Network traffic is composed of packets, as shown in Figure 1, each consisting of two main components: the header (plaintext) and the payload (usually encrypted text). The header contains metadata about the packet, such as source and destination IP addresses, protocol information, and so on. The payload, on the other hand, contains the actual data being transmitted and is often encrypted in the form of hexadecimal to ensure confidentiality and integrity. Current byte-based encrypted traffic classification researches typically use both the header and the payload as inputs to the models. A fixed number of bytes, like 128, 256, etc., are utilized for the classification task. These bytes often include a mix of plaintext information from the header and encrypted data from the payload. At the same time, studies (Akbari et al., 2022; Zhang et al., 2023) have shown that plaintext information significantly contributes to the effectiveness of classification. For example, the paper (Akbari et al., 2022) mentions that the availability of some obvious feature of the server (e.g., plain-text SNI field) in-the-clear is crucial for the utility of deep-learning-based (DL) approaches. Another research (Zhang et al., 2023) proves through ablation experiments that plaintext information is more helpful for the accuracy of classification effect than encrypted text information. This leads to a fascinating hypothesis: Can we achieve satisfactory traffic classification performance using only the plaintext information present in the header?
To further validate the roles of plaintext, encrypted text, and their combination in traffic classification, we compare the one of the state-of-the-art (SOTA) models (Lin et al., 2022; Zhao et al., 2023) performance and time overhead across four datasets. For the combined plaintext-encrypted text setting, as shown in Figure 2(b), the current prevalent methods that utilize both plaintext and encrypted text achieve the highest accuracy (indicated by the blue line). However, this setting results in high time overhead, as depicted in Figure 2(a).
Some studies have demonstrate that the effectiveness of DL-based methods largely rely on plaintext information. This suggests that satisfactory classification performance can be achieved by leveraging only plaintext data. Notably, for the plaintext-only setting, as indicated by the gray line in Figure 2(b), the model achieves around 90% accuracy on certain datasets, further substantiating the crucial role of plaintext in classification. Surprisingly, this setting also enhances time efficiency by approximately 4 to 10 times. However, relying solely on plaintext data obtains lower accuracy on some situations.
For the encrypted text setting, the results in Figure 2(b) indicate that using purely encrypted text for classification performs better than random guessing, implying that encrypted data can also aid in improving classification accuracy alongside plaintext.
In conclusion, the time overhead for the combined plaintext-encrypted text setting is considerably high, although the accuracy is the best in these settings. Therefore, balancing the utilization of plaintext and encrypted text to ensure accurate classification while optimizing time efficiency remains a key research challenge. Our study aims to explore this trade-off to achieve both high accuracy and minimal time overhead in encrypted traffic classification.
To address the above issues, we propose a simple but Efficient and Effective Two-Phase approach (EETP) to balance plaintext and encrypted text for traffic classification. In the first stage, we introduce a new traffic classification task: Determining whether the Plaintext is enough to be accurately Classified (DPC) and propose a DPC selector. The purpose of this stage is to fully leverage plaintext information to quickly and efficiently determine if the traffic can be classified using only plaintext. This allows us to optimize the utilization of time by avoiding unnecessary processing of encrypted data when plaintext alone suffices. In the second stage, we proceed with the plaintext classification model for a efficient classification if the traffic can be classified only using plaintext. Otherwise, a combination of plaintext and encrypted text is considered to enhance the model’s accuracy with the encrypted text classification model. The two-phase design ensures a balance between prediction accuracy and efficiency. By attempting to classify with plaintext, DL- based models can significantly reduce processing time for a large portion of the traffic. For cases where plaintext is insufficient, incorporating encrypted text in the second phase ensures that models still achieve high accuracy as well as low time consumption.
The contributions can be summarized as follows:
  • For the first time, we analyze the impact of plaintext and encrypted text on model performance and time overhead. We introduce an interesting idea of categorizing data packets into two types: plaintext classifiable and plaintext non-classifiable. Based on this categorization, a novel approach is designed to significantly reduce computational overhead while maintaining high prediction accuracy.
  • We propose a two-phase approach to balance the trade-off between plaintext and encrypted text in traffic classification. Stage one is to Determine whether the Plain text is enough to be accurately Classfied (DPC) with the proposed DPC Selector. Stage two aims to adaptively make a classification with the result from stage one.
  • Experimental results on one public dataset and one real-world dataset collected by ourselves demonstrate the efficiency and effectiveness of our approach.
Figure 3. The overview of EETP, which consists of two stages. Stage one is to determine whether the Plain text is enough to be accurately classfied with the proposed DPC Selector. Stage two is to adaptively make a classification with the result from stage one. The green arrows indicate the plaintext processing, brown arrows present the encrypted text processing.
2. Models
The proposed approach EETP is presented in Figure 3, which efficiently and effectively to balance plaintext and encrypted text for traffic classification. The core modules consist of a DPC selector (in stage one) and plaintext or encrypted text classification models (in stage two). Specifically, the DPC selector is able to determining whether the plaintext is enough to be accurately classified. The purpose of this module is to fully leverage plaintext information to quickly and efficiently determine if the traffic can be classified using only plaintext. Then, plaintext or encrypted text classification models can make an adaptively classification with the result from stage one. In the following, we first describe the problem formulation. Then, the detail of the EETP is introduced step by step.
2.1. Problem Formulation
Given a pcap file, the scapy toolkit parses the file and extracts the network packets. Each packet includes header portion (plaintext) $X^{p}$ and payload data (encrypted text) $X^{e}$ in the form of string $X=\{X^{p},X^{e}\}=\{x^{p}_{1},x^{p}_{2},\dots,x^{p}_{M},\dots,x^{e}_{N}\}$ that consists of ${N}$ traffic words, with a category label $y$, e.g. YouTube, Facebook, etc., where $p$ indicates plaintext, $e$ indicates encrypted text, $M$ means the number of plaintext traffic words. Our task can be divided into two subtasks. The first subtask involves determining, based on plaintext information, whether the data is sufficient for accurate classification. This is a binary classification task. The second subtask depends on the result of the first subtask, selecting different branches to output the corresponding packet category label, which is a multi-class classification task.
2.2. Training Stage One
2.2.1. DPC Data Construction
In the first stage of training, the proposed new DPC task involves assessing whether the plaintext information is sufficient for accurate classification. However, existing datasets do not contain labels for this specific purpose, making the construction of the DPC data a crucial step.
Considering that the DPC task is a binary classification problem and relatively straightforward, many SOTA traffic classification models (MengWMLLZ22; Lin et al., 2022) can be leveraged for data labeling. Specifically, we use only plaintext information as input and employ the Pacrep model to perform standard traffic classification training. Once the model is trained, we predict the labels for all data samples. A data sample is labeled as “1” (plaintext classifiable) only if the predicted label matches the original sample’s label. Otherwise, the sample is labeled as “0” (plaintext non-classifiable). The formulation can be defined as follows:
(1) $label_{{\rm{sapmle}}}=\begin{cases}1,&\text{if $y$ == $\hat{y}$}\\ 0,&\text{if $y$ ~{}!= $\hat{y}$}\end{cases}$
where $\hat{y}$ is the predicted label by the traffic classification model.
After obtaining all the labels, we can then train the DPC Selector, which will enable us to distinguish between classifiable and non-classifiable plaintext data effectively. This process ensures accurate and efficient traffic classification in subsequent stages.
Table 1. Comparison results on public and collected datasets. indicates reproduced results. Time metric means that the number of samples that the model predicts per second.
Methods ISCXVPN CHNAPP
Macro-F1 Micro-F1 Time/ each second Macro-F1 Micro-F1 Time/ each second
APPS (Taylor et al., 2018) 81.68 81.33 - - - -
SMT (Shen et al., 2019) 76.31 76.00 - - - -
Deep Packet (Lotfollahi et al., 2020) 65.06 66.70 - - - -
TR-IDS (Min et al., 2018) 54.98 57.66 - - - -
HEDGE (Casino et al., 2019) 29.67 34.00 - - - -
3D-CNN (Zhang et al., 2020) 54.34 57.00 - - - -
BERT (Devlin et al., 2019) 88.45 88.73 - 56.94 84.08 -
PacRep (Meng et al., 2022) 99.32 99.13 22.75 95.10 98.75 44.46
ET-BERT (Lin et al., 2022) 94.25 96.05 - 94.08 96.09 -
YaTC (Zhao et al., 2023) 96.67 98.04 - 90.44 93.50 -
EEPT (Ours) 99.89 99.97 124.01 94.95 99.14 82.11
2.2.2. DPC Selector
After obtaining the training data, we proceed to train the DPC selector. This module aims to filter traffic data, directing plaintext classifiable samples to the plaintext classification model and plaintext non-classifiable samples to the encrypted classification model. This adaptive approach enhances the effectiveness of classification for different types of data, as well as playing a crucial role in optimizing the efficiency of traffic classification. Specifically, existing off-the-shelf traffic classification model  [^1] https://github.com/ict-net/PacRep is utilized as the DPC selector to ensure robust performance. The cross-entropy loss function is used for training. Then, we test the model to selectively separate the samples in the test set into two categories. As illustrated in Figure 3, our experiments show that the majority of data (represented by the green portion) is directed towards the plaintext classification model, while a smaller portion is directed towards the encrypted text classification model. This demonstrates the superiority of our proposed framework, effectively and efficiently achieving traffic classification.
2.3. Training Stage Two
In the second phase, the proposed EETP aims to adaptively select different branches for traffic classification, ensuring both effectiveness and efficiency. Simply, the plaintext classification model and the encrypted classification model utilize the current PLMs-based (Pre-training Language Model) techniques.
3. Experiments
3.1. Public Dataset
First of all, The proposed EETP and comparison methods are evaluated using the public traffic dataset (Draper-Gil et al., 2016). Specifically, the ISCXVPN (Draper-Gil et al., 2016) dataset predominantly consists of network traffic data using Virtual Private Networks (VPNs), with each packet labeled by application, such as Gmail, Facebook, and others.
3.2. Collected Dataset
To further conduct an effective and efficient comparison between the SOTA models and EETP, we conduct a real-world dataset collection manually, named CHNAPP. This dataset comprises network traffic data from five widely-used Internet applications: Weibo, WeChat, QQMail, TaoBao, Youku, and QQMusic. By incorporating these popular applications, CHNAPP provides a robust basis for evaluating and contrasting the performance of EETP with leading models in real-world scenarios. This comprehensive collection ensures that the comparison is both realistic and relevant.
3.3. Evaluation Metrics
The category distribution in these tasks varies, with some being balanced and others imbalanced. To accurately assess performance, both macro F1 (Ma-F1) and micro F1 (Mi-F1) scores are utilized. The macro F1 score calculates the F1 score independently for each class and then takes the average, treating each class equally regardless of its size. This approach is particularly effective for evaluating model performance in scenarios with imbalanced class distributions, as it ensures that smaller classes are given the same consideration as larger ones. Conversely, the micro F1 score aggregates the contributions from all classes to compute a global F1 score. This method places more emphasis on the performance of larger classes, providing a measure of overall accuracy. By considering both macro F1 and micro F1 scores, we can provide a more comprehensive evaluation. Time metric means that the number of samples that the model predicts per second.
3.4. Experimental Setting & Baselines
For the experimental settings, the model parameters are detailed in three aspects. Firstly, for the DPC selector, the training data comprises approximately 100,000 samples, while the testing data consists of 10,000 samples. The training process runs for 3 epochs with a learning rate of 1e-5 and a batch size of 64. Next, for the plaintext classification model, the one-hot layer dimension is set to 100, and the dimension of the 1D-convolution layer is set to (1, 10). This model is trained for 1 epoch with a learning rate of 1e-5 and a batch size of 64. Finally, for the encrypted text classification model, the BERT layer dimension is set to 768. All other parameters remain consistent with the model (Meng et al., 2022). Epoch is set to 5 , with a learning rate of 5e-6 and a batch size of 64. The Adam (Kingma and Ba, 2014) optimizer with the initial learning rate of $1$e-$5$ is utilized for training, and the Pytorch framework (Paszke et al., 2017) is adapted. The experiments are implemented on the Tesla A100-80G GPU.
The baselines can be divided into three categories for comparison: (1) feature-based methods (i.e., APPS (Taylor et al., 2018), SMT (Shen et al., 2019)); (2) byte-based models (i.e., Deep Packet (Lotfollahi et al., 2020), TR-IDS (Min et al., 2018), HEDGE (Casino et al., 2019), 3D-CNN (Zhang et al., 2020)); and (3) PLMs-based methods (i.e., BERT (Devlin et al., 2019), PacRep (Meng et al., 2022), ET-BERT (Lin et al., 2022), YaTC (Zhao et al., 2023)). More details of these compared baselines can be seen in Appendix A.
3.5. Main Results
Considering the different techniques employed by feature-based methods, byte-based models, and PLMs-based methods, we focus our reproduction only on the PLMs-based methods to ensure a fair comparison of time overheads and provide time efficiency metrics for these methods. In addition, to highlight the effectiveness of PLMs-based methods, we compare their macro-F1 and micro-F1 scores with those of feature-based methods and byte-based models. The conclusions drawn from this comparison are as follows: (1) In terms of effectiveness , feature-based methods and byte-based models perform comparably, while PLMs-based methods demonstrate further improvement. Notably, our proposed EEPT model achieves SOTA results across almost all evaluation metrics, illustrating the superiority of the two-stage framework, which adaptively make a balance between the plaintext and encrypted text. (2) From a efficiency perspective, EEPT exhibits a 2 to 5 times improvement across two datasets. Specifically, when compared to the pacrep model—a BERT-based model fine-tuned directly under supervised learning—the EEPT model shows clear advantages. ET-BERT and YaTC, on the other hand, incorporate an additional pre-training phase before fine-tuning, resulting in greater overhead compared to directly fine-tuning BERT. Overall, the analysis of both effectiveness and efficiency demonstrates the superiority of the EEPT framework and underscores the necessity of analyzing both plaintext and encrypted text.
4. Conclusion
In this paper, we propose a simple but Efficient and Effective Two-Phase approach (EETP) to balance plaintext and encrypted text for traffic classification. Stage one is to determine whether the plain text information is enough to be accurately classified using the proposed DPC Selector. Stage two aims to adaptively make a classification with the result from stage one. The proposed two-phase framework provides a new insight to make an analysis on the plaintext and encrypted text in network traffic field. In summary, EETP is very effective and efficient, which outperforms the previous methods on two datasets. For future work, the design of modules in the EEPT could be considered for further research.
Appendix A Details of Baselines
The details of utilized baselines can be seen in the following.
  • APPS (Taylor et al., 2018), which selects statistical features, and uses a random tree classifier for application classification.
  • SMT (Shen et al., 2019), which fuses features of different dimensions by a kernel function for decentralized applications fingerprinting.
  • Deep Packet (Lotfollahi et al., 2020), which integrates both feature extraction and classification phases to train the stacked autoencoder and CNN for application classification.
  • 3D-CNN (Zhang et al., 2020), which leverages the byte data as input for 3D-CNN and can classify packets from both known and unknown patterns.
  • BERT (Devlin et al., 2019), which is based on Transformer architecture and has achieved noticeable performance on various NLP tasks.
  • PacRep (Meng et al., 2022), which utilizes the BERT as an encoder and uses contrastive loss to optimize the learned packet representations with multi-task learning.
  • ET-BERT (Lin et al., 2022), which pre-trains deep contextualized datagram-level representation from large-scale raw traffic bytes.
  • YaTC (Zhao et al., 2023), which introduces a masked autoencoder based traffic transformer with multi-level flow representation and then performs pre-training process for traffic classification.